# Measuring Ransomware Lateral Movement Susceptibility via Privilege-Weighted Adjacency Matrix Exponentiation Satyam Tyagi1 Ganesh Murugesan1

### v1

@@
Abstract: Ransomware impact hinges on how easily an intruder can move laterally and spread to the maximum number of assets. We present a graph-theoretic method to measure lateral-movement susceptibility and estimate blast radius. We build a directed multigraph where vertices represent assets and edges represent reachable services (e.g., RDP/SSH) between them. We model lateral movement as a probabilistic process using a pivot potential factor $\pi(s)$ for each service. This allows us to iteratively compute a $K$-hop compromise probability matrix that captures how compromise propagates through the network. Metrics derived from this model include: (1) Lateral-Movement Susceptibility (LMSK): the average probability of a successful lateral movement between any two assets (0-1 scale); and (2) Blast-Radius Estimate (BREK): the expected percentage of assets compromised in an average attack scenario. Interactive control (SSH 22, RDP 3389) gets higher $\pi(s)$ than app-only ports (MySQL 3306, MSSQL 1433), which seldom enable pivoting without an RCE. Across anonymized enterprise snapshots, pruning high-$\pi(s)$ edges yields the largest LMSK/BREK drop, aligning with CISA guidance, MITRE ATT&CK (TA0008: Lateral Movement), and NIST SP 800-207. The framework evaluates (micro)segmentation and helps prioritize controls that reduce lateral movement susceptibility and shrink blast radius.
Keywords: adjacency matrix; multi-hop reachability; privilege weighting; lateral movement; microsegmentation; blast radius; ransomware; Zero Trust
@@ 1.  Introduction
We introduce a privilege- and account-aware adjacency model that composes per-service “pivot potentials” $\pi(s)\!\in\![0,1]$ with multiplicative step success and a noisy-OR union across alternative routes. Iterating this algebraic path closure yields (i) a $k$-hop (or $\varepsilon$-fixed-point) compromise probability matrix and (ii) two deployment metrics: LMSK (lateral-movement susceptibility; mean off-diagonal) and BREK (blast-radius estimate; mean including the diagonal).
@@
Once an adversary holds valid credentials and an interactive remote service is exposed (RDP/SSH), lateral movement does not require unpatched vulnerabilities: adversaries log into remote services and act as the authenticated user [MITRE, 2025f, b, c]. Patch status is orthogonal to pivotability when privileged landing is possible; hence our operators weight service exposure and privilege class, not CVE likelihoods.
@@
  • Privilege- and account-aware path closure. We formalize lateral movement over $[0,1]$ with step-wise multiplication and noisy-OR unions, explicitly capturing valid-account pivots on interactive services [MITRE, 2025b].
@@
  • Theoretical bounds. We situate our iteration within algebraic path theory, giving a series upper bound via $U_{K}=A+AW+\cdots+AW^{K-1}$ and a single-path product lower bound.
  • Alignment with guidance. Our empirical reductions in LMSK/BREK under RDP/SSH constraints match Zero Trust and ransomware guidance to limit remote desktop services and constrain lateral movement [Rose et al., 2020; CISA, 2025a, b, c].
@@
Zero Trust (NIST SP 800-207) calls for minimizing implicit trust and continuously evaluating access in context [Rose et al., 2020]. Our metrics quantify the residual ease of movement under candidate policies, complementing microsegmentation by highlighting edges whose removal yields the steepest probability-mass drop.
@@
Section 2 reviews the literature. Section 3 formalizes matrices, operators, and fixed-point computation. Section 4 defines LMSK/BREK. Section 5 presents properties and bounds. Section 6 provides examples and evidence.
@@ 2.  Background and Related Work
Account-based pivots vs. vulnerability-driven traversal. Beyond exploit-based movement (e.g., T1210) [MITRE, 2025d], ATT&CK catalogs techniques enabling lateral movement without software flaws: Valid Accounts (T1078) and Remote Services (T1021, including RDP/SSH) describe abusing legitimate credentials to authenticate and operate on remote hosts [MITRE, 2025f, b, c, e]. Our formulation treats interactive services as high-pivot edges when paired with account access, independent of CVE posture.
@@
Attack graphs and probabilistic security metrics. Classical attack graphs model preconditions/exploits and enumerate paths to goals [Phillips and Swiler, 1998; Sheyner et al., 2002]. Later work augments graphs with probabilities and costs. We deliberately do not encode per-CVE logic; instead we treat reachable services as coarse pivot opportunities whose probabilities compose multiplicatively with noisy-OR unions, trading granularity for deployability on routine network snapshots.
@@
Algebraic path problems and semiring views. Many path computations admit uniform treatment via algebraic path frameworks replacing $(+,\times)$ by problem-specific $(\oplus,\otimes)$. We adopt multiplication for per-step success and probabilistic sum (noisy-OR) for unions, yielding a monotone iteration toward a least fixed point; this aligns with semiring-style closures, though our $\oplus$ is non-idempotent [Mohri, 2002; Sanmartín and Naranjo, 2022].
@@
Noisy-OR and causal independence. Our union operator is the standard noisy-OR from probabilistic graphical models, aggregating independent causal mechanisms with low elicitation burden [Heckerman, 1994]. Generalizations (e.g., Noisy-MAX) clarify when such factorizations are appropriate.
@@
Epidemic spreading analogies. Multiplicative transmission along edges with union across parallel paths is mathematically analogous to network contagion; classic results connect topology to spread [Newman, 2002]. Our aim is risk measurement (expected reach/blast), not epidemic prediction, but the analogy explains the outsized effect of a small set of high-pivot edges.
@@
AD attack-path tooling. In enterprise Windows domains, BloodHound operationalizes graph-based lateral-movement analysis at scale [SpecterOps, 2025]. We follow similar instincts (directed, typed edges) but compute a probability-weighted closure and provide system-level metrics for control evaluation.
@@
Zero Trust and ransomware guidance. NIST SP 800-207 sets the architectural principle; ransomware advisories emphasize limiting RDP/remote services to reduce both initial access and lateral movement [Rose et al., 2020; CISA, 2025a, b, c].
2.1 Graph-theoretic underpinnings
$k$-hop walks via adjacency powers.
Let $A$ be the (directed) adjacency matrix of the network. Classical algebraic graph theory shows that $(A^{k})_{ij}$ equals the number of $i{\to}j$ walks of length kk. [^1] As usual, walks may repeat vertices; paths are walks with no repeated vertices. Our probabilistic closure works with walks. See, e.g., Godsil and Royle [2001]; Biggs [1993]; Diestel [2005]. This formalizes the “$k$-hop” notion we use when bounding series $A+AW+\cdots+AW^{K-1}$.
Boolean-OR reachability and transitive closure.
Over the Boolean semiring $(\lor,\land)$, the transitive closure (reachability) can be computed by Boolean matrix powering and union,
$R\;=\;A\lor A^{(2)}\lor\cdots\lor A^{(n)}\,$,
or equivalently via Warshall’s dynamic program for transitive closure [Warshall, 1962]. From a more algebraic perspective, Kleene-style closure over idempotent semirings and dioids gives a unifying view of reachability and shortest paths [Kozen, 1994; Gondran and Minoux, 2008].
Noisy-OR style union for probabilistic graphs.
@@ 2.1 Graph-theoretic underpinnings
When parallel, conditionally independent mechanisms can each enable a step (e.g., multiple services/credentials offering a pivot), the probability that at least one succeeds is the probabilistic sum
@@
the same aggregation used in network reliability for parallel components [Colbourn, 1987]. In probabilistic graphical models, this is the standard noisy-OR (and its generalizations), which captures causal independence with one parameter per parent [Pearl, 1988; Heckerman, 1994; Diez, 1993]. Our use of a noisy-OR union along with multiplicative step success places the model within these classical frameworks.
@@ 3.  The Proposed Model
Our model quantifies lateral movement exposure by representing the network as a directed graph and simulating the propagation of compromise probability. This requires two foundational matrices derived from network state and security policies.
3.1 Foundational Matrices
  1. 1.
    The Adjacency Matrix ($A$): This is an $n\times n$ binary matrix representing the network’s underlying structure, where $n$ is the number of assets. An edge exists only if a service is both permitted by network policy and actively listening on the destination asset.
[…]
    We take $A_{ii}=0$ for all $i$ (no self-loops). Let $I$ denote the $n\times n$ identity matrix.
  2. 2.
    The Pivot Potential Matrix ($W$): This is an $n\times n$ matrix that encodes the inherent pivot potential of each edge. Instead of considering only the single most powerful service, the model treats each reachable service on an edge as an independent opportunity for an attacker. The entry $(W)_{ij}$ represents the total probability of a successful pivot from asset $i$ to $j$, calculated as the probabilistic union (or ”noisy-OR”) of the pivot potentials of all services available on that edge.
    Let $S_{ij}$ be the set of all services reachable from asset $i$ to asset $j$. The pivot potential is calculated as:
    $W_{ij}=1-\prod_{s\in S_{ij}}(1-\pi(s))$
    Convention. $S_{ij}$ includes only services that are (i) allowed by policy and (ii) observed listening on $j$; hence $W_{ij}>0\Rightarrow A_{ij}=1$ (i.e., $W\leq A$ elementwise). In particular, if $S_{ij}=\varnothing$ then $W_{ij}=0$.
3.2 Computing Reachability (Unweighted Method)
Before detailing our probabilistic model, we first describe the standard, unweighted method for calculating K-hop reachability. This method answers the binary question: ”Does a path of length $K$ or less exist?” This is computed iteratively. Let $S_{k}$ be the matrix where $(S_{k})_{ij}=1$ if a path of at most $k$ hops exists.
  • Base Case: The 1-hop reachability matrix is the adjacency matrix, $S_{1}=A$.
  • Iteration: The reachability matrix for at most $k$ hops, $S_{k}$, is found by taking the previous state $S_{k-1}$ and adding any new paths formed by one additional hop:
    $S_{k}=S_{k-1}\lor(S_{k-1}\cdot_{bool}A)$
  • Final Result: The result is the matrix $S_{K}$ after K1K-1 iterations.
While this method confirms path existence, it treats a high-risk path (e.g., via SSH) as identical to a low-risk one (e.g., via a database port), which our probabilistic model addresses. Let $S_{1}=I\lor A$ and $S_{k+1}=S_{k}\lor(S_{k}A)$; write $S_{K}$ for the $K$-hop boolean reachability matrix.
@@ 3.3 Computing Compromise Probability (Probabilistic Method)
    The Probabilistic Composition Product ($\otimes$): This operator, replacing matrix multiplication, calculates the probability of new compromise paths forming by propagating existing compromise probabilities through the pivot potentials in $W$. For matrices $X$ and $Y$, the product $Z=X\otimes Y$ is:
    $Z_{ij}=(X\otimes Y)_{ij}=1-\prod_{m=1}^{n}(1-X_{im}\times Y_{mj})$ (1)
@@
    The Probabilistic Union Operator ($\lor_{p}$): This operator, replacing matrix addition, correctly combines probabilities from independent sets of paths using the union rule. For matrices $X$ and $Y$, the union $Z=X\lor_{p}Y$ is:
    $Z_{ij}=(X\lor_{p}Y)_{ij}=1-(1-X_{ij})(1-Y_{ij})$ (2)
@@
It is important to note that the probabilistic composition operator ($\otimes$) implicitly treats paths through different intermediate assets ($m$) as probabilistically independent events. This is a standard assumption in many network models that makes the problem computationally tractable. We acknowledge that in some real-world scenarios, such as an attack using a single stolen credential, these paths may not be fully independent.
Conservative bound. For correlated routes, we can replace $1-\prod_{m}(1-X_{im}Y_{mj})$ with $\max_{m}(X_{im}Y_{mj})$ (lower bound).
@@
The 1-hop compromise probability matrix, $P_{1}$, is the adjacency matrix $A$ with a diagonal of 1s, as an asset that is the starting point of an attack is by definition compromised.
@@
For each hop from $k=1$ to K1K-1, we calculate the next state, $P_{k+1}$. The new state is the probabilistic union of the current state, $P_{k}$, with all the new compromise paths that can be formed by one additional hop.
$P_{k+1}=P_{k}\lor_{p}(P_{k}\otimes W)$
@@
After K1K-1 iterations, the resulting matrix is the final K-hop Compromise Probability Matrix, $P_{K}$.
Hop-1 semantics. We treat the first hop as pure network reachability: $P_{1}=I\lor_{p}A$. Pivot potentials in $W$ affect only subsequent hops ($k\geq 2$) via $P_{k}\otimes W$.

### v2

@@
Abstract: Ransomware impact hinges on how easily an intruder can move laterally and spread to the maximum number of assets. We present a graph-theoretic formulation that casts lateral movement as a path-closure problem over a probability semiring to measure lateral-movement susceptibility and estimate blast radius. We build a directed multigraph where vertices represent assets and edges represent reachable services (e.g., RDP/SSH) between them. We model lateral movement as a probabilistic process using a pivot potential factor $\pi(s)$ for each service, with step successes composed via a probabilistic path operator $\otimes$ and alternative paths aggregated via a probabilistic union $\oplus$ (noisy-OR). This yields a monotone fixed-point (iterative) computation of a $K$-hop compromise probability matrix that captures how compromise propagates through the network. Metrics derived from this model include: (1) Lateral-Movement Susceptibility (LMSK): the average probability of a successful lateral movement between any two assets (0-1 scale); and (2) Blast-Radius Estimate (BREK): the expected percentage of assets compromised in an average attack scenario. Interactive services (SSH 22, RDP 3389) receive higher $\pi(s)$ than app-only ports (MySQL 3306, MSSQL 1433), which seldom enable pivoting without an RCE. Across anonymized enterprise snapshots, pruning high-$\pi(s)$ edges yields the largest LMSK/BREK drop, aligning with CISA guidance, MITRE ATT&CK (TA0008: Lateral Movement), and NIST SP 800-207. The framework evaluates (micro)segmentation and helps prioritize controls that reduce lateral-movement susceptibility and shrink blast radius.
Keywords: adjacency matrix; multi-hop reachability; privilege weighting; probability semiring; fixed-point iteration; lateral movement; microsegmentation; blast radius; ransomware; Zero Trust
@@ 1.  Introduction
We introduce a privilege- and account-aware adjacency model, casting lateral movement as a path-closure problem over a probability semiring. The model composes per-service “pivot potentials” $\pi(s)\!\in\![0,1]$ with a probabilistic path operator $\otimes$ for sequential step composition and a probabilistic union $\oplus$ across alternative routes (noisy-OR). Iterating this algebraic path-closure yields (i) a $K$-hop (or $\varepsilon$-fixed-point) compromise probability matrix and (ii) two deployment metrics: LMSK (lateral-movement susceptibility; mean off-diagonal) and BREK (blast-radius estimate; mean including the diagonal).
@@
Once an adversary holds valid credentials and an interactive remote service is exposed (RDP/SSH), lateral movement does not require unpatched vulnerabilities: adversaries log into remote services and act as the authenticated user [6, 2, 3]. Patch status is orthogonal to pivotability when privileged landing is possible; hence our operators weight service exposure and privilege class, not CVE likelihoods.
@@
  • Privilege- and account-aware path closure. We formalize lateral movement over $[0,1]$ with probabilistic path composition ($\otimes$) and probabilistic union ($\oplus$), explicitly capturing valid-account pivots on interactive services [2].
@@
  • Theoretical bounds. We show that our probabilistic model $P_{K}$ is upper-bounded (entrywise) by the $K$-hop Boolean reachability $R_{K}$ and establish a single-path product lower bound (cf. Section 6).
  • Alignment with guidance. Our empirical reductions in LMSK/BREK under RDP/SSH constraints match Zero Trust and ransomware guidance to limit remote desktop services and constrain lateral movement [7, 8, 9, 10].
@@
Zero Trust (NIST SP 800-207) calls for minimizing implicit trust and continuously evaluating access in context [7]. Our metrics quantify the residual ease of movement under candidate policies, complementing microsegmentation by highlighting edges whose removal yields the steepest probability-mass drop.
@@
Section 2 reviews the literature; Section 3 develops the graph-theoretic underpinnings; Section 4 formalizes the matrices, operators ($\oplus$ and $\otimes$), and computation for the proposed model; Section 5 defines the metrics LMSK and BREK; Section 6 presents formal properties and bounds; Section 7 provides examples and evidence; and Section 8 concludes with discussion and future work.
@@ 2.  Background and Related Work
Account-based pivots vs. vulnerability-driven traversal. Beyond exploit-based movement (e.g., T1210) [4], ATT&CK catalogs techniques enabling lateral movement without software flaws: Valid Accounts (T1078) and Remote Services (T1021, including RDP/SSH) describe abusing legitimate credentials to authenticate and operate on remote hosts [6, 2, 3, 5]. Our formulation treats interactive services as high-pivot edges when paired with account access, independent of CVE posture.
@@
Attack graphs and probabilistic security metrics. Classical attack graphs model preconditions/exploits and enumerate paths to goals [11, 12]. Later work augments graphs with probabilities and costs. We deliberately do not encode per-CVE logic; instead we treat reachable services as coarse pivot opportunities whose probabilities compose via a probabilistic path operator ($\otimes$) with probabilistic union ($\oplus$) across alternatives, trading granularity for deployability on routine network snapshots.
@@
Algebraic path problems and operator view. Many path computations admit uniform treatment via algebraic path frameworks replacing $(+,\times)$ by problem-specific $(\oplus,\otimes)$. We adopt a probabilistic path operator $\otimes$ for per-step composition and a probabilistic union $\oplus$ for alternatives (noisy-OR), yielding a monotone iteration toward a least fixed point; this is analogous to algebraic path problems—we do not assume a semiring ($\oplus$ and $\otimes$ are non-idempotent, and distributivity holds only as an inequality) [13, 17].
@@
Probabilistic union (noisy-OR) and causal independence. Our probabilistic union operator corresponds to the standard noisy-OR from probabilistic graphical models, aggregating independent causal mechanisms with low elicitation burden [14]. Generalizations (e.g., Noisy-MAX) clarify when such factorizations are appropriate.
@@
Epidemic spreading analogies. Probabilistic composition along edges with probabilistic union across parallel paths is mathematically analogous to network contagion; classic results connect topology to spread [15]. Our aim is risk measurement (expected reach/blast), not epidemic prediction, but the analogy explains the outsized effect of a small set of high-pivot edges.
@@
AD attack-path tooling. In enterprise Windows domains, BloodHound operationalizes graph-based lateral-movement analysis at scale [16]. We follow similar instincts (directed, typed edges) but compute a probability-weighted closure and provide system-level metrics for control evaluation.
@@
Zero Trust and ransomware guidance. NIST SP 800-207 sets the architectural principle; ransomware advisories emphasize limiting RDP/remote services to reduce both initial access and lateral movement [7, 8, 9, 10].
3.  Graph-theoretic underpinnings
3.1 Adjacency matrices and notation
Let $A$ be the (directed) adjacency matrix of the network; $A_{ij}{=}1$ iff a policy-permitted, actively listening service on $j$ is reachable from $i$, and $A_{ii}{=}0$; later we use a pivot-potential matrix $W\!\in[0,1]^{n\times n}$ with $W\leq A$ (elementwise).
[…]
$A=\begin{bmatrix}0&1&1&0\\ 0&0&1&1\\ 0&0&0&1\\ 0&0&0&0\end{bmatrix}$
Figure 1: Directed graph and its adjacency matrix AA.
3.2 $K$-hop path reachability (Boolean closure)
Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix ($A_{ij}{=}1$ iff there is an edge $i{\to}j$). Over the Boolean semiring $(\lor,\land)$, define the Boolean matrix product $(X\diamond Y)_{ij}:=\bigvee_{\ell=1}^{n}X_{i\ell}\land Y_{\ell j}$, Boolean powers $A^{[0]}:=I$, $A^{[1]}:=A$, and for $k\geq 1$, $A^{[k{+}1]}:=A^{[k]}\diamond A$.
Claim (standard). For any $i,j$ and $k\geq 1$, $A^{[k]}_{ij}=1$ iff there exists a simple path from $i$ to $j$ of length at most $k$. (Proof sketch: if a $k$-step walk exists, prune cycles to obtain a simple path of length $\leq k$; conversely, a $k$-edge path witnesses $A^{[k]}_{ij}=1$.) See [27, §1.3].
Hence the $K$-bounded path reachability is
$R_{K}\;=\;I\lor A\lor A^{[2]}\lor\cdots\lor A^{[K]},$
and the full transitive closure is $R_{n-1}$, since any walk of length $\geq n$ repeats a vertex and can be shortened without changing reachability. One may compute $R_{n-1}$ via Warshall’s dynamic program or view it as a Kleene-style closure over an idempotent semiring [20, 21, 22].
[…]
$A^{[2]}\;=\;A^{2}\ \text{(Boolean)}\;=\;\begin{bmatrix}0&0&1&1\\ 0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix}$
Figure 2: Length-2 paths. Boolean square $A^{[2]}$ (entry $(u,v)=1$ iff some $w$ gives $u\!\to\!w\!\to\!v$ under Boolean product).
[…]
$R^{(\leq 2)}\;=\;I\,\lor\,A\,\lor\,A^{[2]}\;=\;\begin{bmatrix}1&1&1&1\\ 0&1&1&1\\ 0&0&1&1\\ 0&0&0&1\end{bmatrix}$
Figure 3: Boolean-OR reachability up to length 2. Combines self-reachability, 1-step, and 2-step paths.
Computing reachability in practice (unweighted).
We compute the $K$-bounded Boolean-OR reachability recursively using the Boolean product $\diamond$ and union $\lor$.
  • Base case. $R_{1}\;=\;I\,\lor\,A.$
  • Iteration. For $k\geq 1$,
    $R_{k+1}\;=\;R_{k}\,\lor\,(R_{k}\diamond A),$
    i.e., add all new (k+1)(k{+}1)-hop connections obtainable by one more Boolean step from the current frontier $R_{k}$.
  • Final result / stopping. After K1K{-}1 iterations, return $R_{K}$. Equivalently, stop early at the fixed point if $R_{k+1}=R_{k}$ (no new reachability).
Notes. This iterative union of Boolean powers is equivalent to Warshall’s transitive-closure algorithm and to the Kleene star on the Boolean semiring. In practice, bitset/word-parallel Boolean matrix multiplication for $\diamond$ makes each iteration fast; complexity is comparable to $O(n^{3})$ in dense form, with early stopping common in sparse graphs.
3.3 Pivot potential
Definition and example.
Let $S_{ij}$ denote the enabling mechanisms (e.g., services/credentials) from $i$ to $j$. The pivot-potential matrix $W\in[0,1]^{n\times n}$ aggregates mechanisms on the same edge via probabilistic union, i.e.
$W_{ij}\;=\;1-\prod_{m\in S_{ij}}\bigl(1-\pi(m)\bigr),\qquad W\leq A\text{ (elementwise)}.$
[…]
$W=\begin{bmatrix}0.00&0.80&0.30&0.00\\ 0.00&0.00&0.80&0.80\\ 0.00&0.00&0.00&0.30\\ 0.00&0.00&0.00&0.00\end{bmatrix}$
Figure 4: Pivot potential matrix $W$ via noisy-OR aggregation over services on each edge.
Probabilistic union (noisy-OR) across mechanisms.
@@ 3.3 Pivot potential
When parallel, conditionally independent mechanisms can each enable a step (e.g., multiple services/credentials offering a pivot), the probability that at least one succeeds is the probabilistic union
@@
the same aggregation used in network reliability for parallel components [23]. In probabilistic graphical models, this is the standard noisy-OR (and its generalizations), which captures causal independence with one parameter per parent [24, 14, 25]. In our framework we denote this union by $\oplus$ and pair it with a probabilistic path composition $\otimes$ for sequential steps, placing the model within classical algebraic-path/semiring views.
@@ 4.  The Proposed Model
Our model quantifies lateral-movement exposure on the directed graph using the matrices $A$ and $W$ defined in Sec. 3, and computes $K$-hop compromise probabilities. We now describe the probabilistic operators and iteration.
@@ 4.1 Computing Compromise Probability (Probabilistic Method)
    The Probabilistic Composition Product ($\otimes$): This operator, replacing matrix multiplication, composes sequential steps and aggregates alternatives via the probabilistic union. For matrices $X$ and $Y$,
    $Z_{ij}\;=\;(X\otimes Y)_{ij}\;=\;\bigoplus_{m=1}^{n}\bigl(X_{im}\,\otimes\,Y_{mj}\bigr)\;=\;1-\prod_{m=1}^{n}\Bigl(1-\bigl(X_{im}\,\otimes\,Y_{mj}\bigr)\Bigr).$ (1)
@@
    The Probabilistic Union Operator ($\oplus$): This operator, replacing matrix addition, correctly combines probabilities from independent sets of paths using the union rule. For matrices $X$ and $Y$, the union $Z=X\oplus Y$ is:
    $Z_{ij}=(X\oplus Y)_{ij}=1-(1-X_{ij})(1-Y_{ij})$ (2)
@@
It is important to note that the probabilistic composition ($\otimes$) treats the success events through different intermediates $m$ as conditionally independent. This is a standard assumption in many network models that makes the problem computationally tractable. We acknowledge that in some real-world scenarios, such as an attack using a single stolen credential, these events may not be fully independent.
Conservative bound. For conditionally dependent routes, replace  $1-\prod_{m}\!\bigl(1-(X_{im}\otimes Y_{mj})\bigr)\,$ with $\max_{m}\bigl(X_{im}\otimes Y_{mj}\bigr)$ (lower bound).
@@
The 1-hop compromise probability matrix uses edge step-success on the first hop:
$P_{1}\;=\;I\,\oplus\,W.$
[…]
$P_{1}\;=\;I\,\oplus\,W\;=\;\begin{bmatrix}1.00&0.80&0.30&0.00\\ 0.00&1.00&0.80&0.80\\ 0.00&0.00&1.00&0.30\\ 0.00&0.00&0.00&1.00\end{bmatrix}$
Figure 5: Hop-1 baseline. $P_{1}=I\oplus W$ encodes self-reachability and probabilistic 1-hop step-success on edges.
@@
For each hop from $k=1$ to K1K-1, let $W_{1}:=W$ and define $W_{k+1}=W_{k}\otimes W$. Then compute
$P_{k+1}\;=\;P_{k}\,\oplus\,W_{k+1}.$
Note. Since $W_{1}=W$ already contains all 1-hop terms and $P_{k}$ accumulates $I\oplus W_{1}\oplus\cdots\oplus W_{k}$, the union with $P_{k}$ leaves shorter-hop mass unchanged and only adds longer-hop contributions.
[…]
$W_{2}\;=\;W_{1}\otimes W\;=\;\begin{bmatrix}0.00&0.00&0.64&0.67\\ 0.00&0.00&0.00&0.24\\ 0.00&0.00&0.00&0.00\\ 0.00&0.00&0.00&0.00\end{bmatrix}$
Figure 6: Two-hop contribution. $W_{2}=W\otimes W$ shows exactly 2-hop mass.
@@
After K1K{-}1 iterations, return the cumulative matrix
$P_{K}\;=\;P_{K-1}\oplus W_{K}\;=\;I\oplus W_{1}\oplus W_{2}\oplus\cdots\oplus W_{K},\quad\text{with }W_{1}:=W,\;W_{k+1}=W_{k}\otimes W.$
[…]
$P_{2}\;=\;P_{1}\oplus W_{2}\;=\;\begin{bmatrix}1.00&0.80&0.75&0.67\\ 0.00&1.00&0.80&0.85\\ 0.00&0.00&1.00&0.30\\ 0.00&0.00&0.00&1.00\end{bmatrix}$
Figure 7: Up-to-2-hop compromise probability. Combine hop-1 with the 2-hop contribution: $P_{2}=P_{1}\oplus W_{2}$.